A machinist turns on the power on to a grinding wheel at time t = 0 s. The wheel accelerates uniformly from rest for 10 s and reaches the operating angular speed of 96 rad/s The wheel is run at that angular velocity for 40 s and then power is shut off. The wheel slows down uniformly at 1.5 rad/s^2 until the wheel stops. For how long a time after the power is shut off does it take the wheel to stop?

A) 64 s

B) 62 s

C) 66 s

D) 68 s

E) 70 s
Answer #1

angular acceleration = alpha = - 1.5 rad/sec^2 ( - ve is slows down )

initial angular velocity = w0 = 96 rad/sec

when it stops angular velocity = w = 0

               W = W0 + alpha * t.

          time = t = w- w0 / alpha

                   t = 96 / 1.5 = 64 sec.................. A